### v1

# Multiple shells in supernova 2023ixf support the jittering jets explosion mechanism (JJEM)

Abstract
Examining the photospheric time evolution of the core-collapse supernova (CCSN) SN 2023ixf from the literature, we identify three evolutionary time periods with constant expansion velocities, which we attribute to three ejecta shells. We find that several CCSN remnants have morphologies with two or more complete or partial shells, compatible with the presence of two or three photospheric shells during the photospheric phase of the explosion. Studies have attributed these CCSN remnants to the jittering-jet explosion mechanism (JJEM), which involves two or three energetic pairs of jets participating in the explosion. We, therefore, conclude that the structure of the photospheric shells of SN 2023ixf supports its explosion by the JJEM. This study adds to the accumulating evidence that the JJEM is the primary explosion mechanism of CCSNE.
Supernova remnants – Massive stars – Circumstellar material – Stellar jets – Supernova: individual (SN 2023ixf)
1 Introduction
A major question about core-collapse supernovae (CCSNe) is the mechanism of their explosion. There are two intensively studied competing theoretical explosion mechanisms of core-collapse supernovae (CCSNe), the delayed neutrino mechanism and the jittering jets explosion mechanism (JJEM).
The research on the delayed neutrino explosion mechanism focuses on simulating the revival of the stalled shock around the newly-born neutron star (NS) at $r\simeq 150~\rm{km}$ by neutrino heating, finding the conditions for an explosion rather than a ‘failed supernova’, and comparing simulated outcomes with observations (e.g., Bamba et al. 2025; Boccioli et al. 2025; Boccioli & Roberti 2025; Eggenberger Andersen et al. 2025; Fang et al. 2025; Huang et al. 2025; Imasheva et al. 2025; Laplace et al. 2025b; Maltsev et al. 2025; Maunder et al. 2024; Mori et al. 2025; Müller et al. 2025; Nakamura et al. 2025; Sykes & Müller 2025; Janka 2025a; Orlando et al. 2025; Paradiso & Coughlin 2025; Tsuna et al. 2025; Vink et al. 2025; Wang & Burrows 2025; Willcox et al. 2025; Mukazhanov 2025; Raffelt et al. 2025; Vartanyan et al. 2025). The magnetorotational explosion mechanism operates in very rare cases of rapid pre-collapse core rotation, leading to the launch of one pair of jets along a fixed axis (e.g., Shibata et al. 2025). At best, it can account for only a small fraction of CCSNe, and studies of the magnetorotational explosion mechanism attribute most CCSNe to the neutrino-driven mechanism; therefore, we group the magnetorotational mechanism with the neutrino-driven mechanism. [^1] See Janka 2025b for a recent talk on the neutrino driven mechanism: https://www.memsait.it/videomemorie/volume-2-2025/VIDEOMEM_2_2025.46.mp4; also https://www.youtube.com/watch?v=nRfDPPSmnzI&t=100s
The research of the JJEM has focused since 2024 on finding supporting evidence for jets in CCSN remnants (CCSNRs; Bear et al. 2025; Bear & Soker 2025; Shishkin et al. 2025; Soker 2025a, b, c, d; Soker & Akashi 2025; Soker & Shishkin 2025a, b, for papers since 2025) [^2] See Soker 2025e For a talk on the JJEM: https://www.memsait.it/videomemorie/volume-2-2025/VIDEOMEM_2_2025.47.mp4. The morphologies of many CCSNRs, in particular point-symmetric CCSNRs (those with two or more symmetry axes), strongly support the JJEM (see the summary of results in, e.g., Soker 2024a, 2025f). Other studies of the JJEM since 2025 are three-dimensional simulations of jet shaping (Braudo et al., 2025) and pre-collapse angular momentum fluctuations in the inner core (Wang et al., 2025).
The explosion process of CCSNe provides only a few observables to distinguish between the two mechanisms. A minority of CCSNe that have explosion energies of $E_{\rm ex}\gtrsim 2\times 10^{51}~\rm{erg}$ support the JJEM because the neutrino-driven mechanism struggles to reach these explosion energies (for recent reviews see Soker 2024a, 2025f). This is the case, for example, in many superluminous stripped-envelope supernovae (e.g., Kumar 2025). However, the two explosion mechanisms have similar predictions to most other observables of the explosion process itself (before the remnant is spatially resolved; Soker 2024a, 2025f). In this study, we find that SN 2023ixf has a property that is more compatible with the JJEM.
We use the photosphere radius evolution $R_{\rm ph}(t)$ of SN 2023ixf that Zimmerman et al. (2024) calculated (Section 2) to claim that jets exploded SN 2023ixf. In Section 3, we present CCSNRs that have jet-shaped morphologies that are compatible with two or more photospheric shells (depending on the line of sight) as we deduced in Section 2. In this study, shells are not necessarily spherical; they can be caps of protrusions termed ears or partial shells, such as the dense boundary of a jet-inflated bubble. We summarize this short study in Section 4.
2 Identifying photospheric shells in SN 2023ixf
SN 2023ixf is a well-studied nearby CCSN (e.g., Soraisam et al. 2023; Bostroem et al. 2024, 2025; Van Dyk et al. 2024a; DerKacy et al. 2025; Jacobson-Galán 2025; Jacobson-Galán et al. 2025; Medler et al. 2025; Park et al. 2025; Vinko et al. 2025; Zheng et al. 2025). It attracted attention from the beginning because of its proximity and dense, compact circumstellar material (e.g., Berger et al. 2023; Bostroem et al. 2023; Grefenstette et al. 2023; Jacobson-Galán et al. 2023; Smith et al. 2023; Teja et al. 2023; Laplace et al. 2025a). There are uncertainties regarding the progenitor and explosion energy. Singh et al. (2024) model the lightcurve with hydrodynamical calculations and estimated a progenitor radius of $R_{\ast}\simeq 470R_{\odot}$ and an explosion energy of $E_{\rm exp}\simeq 2\times 10^{51}~\rm{erg}$. However, they commented that these values are not unique due to modeling degeneracies. Van Dyk et al. 2024b estimated the progenitor stellar radius to be in the range of $\simeq 921-2012R_{\odot}$, with a median of $R_{\ast}=1389R_{\odot}=9.7\times 10^{13}~\rm{cm}$. Hsu et al. (2025) estimated the progenitor radius to be $R_{\ast}\gtrsim 950R_{\odot}$. Moriya & Singh (2024) estimated the explosion energy to be $E_{\rm exp}\simeq(2-3)\times 10^{51}~\rm{erg}$, while Michel et al. (2025) estimated a more typical energy of $E_{\rm exp}\simeq(0.3-1.4)\times 10^{51}~\rm{erg}$.
The parameters we take to our study are a typical CCSN explosion energy, and a progenitor radius in the range of $R_{\ast}\simeq 600R_{\odot}-1400R_{\odot}$, or $R_{\ast}\simeq 4\times 10^{13}-10^{14}~\rm{cm}$.
We examine the photospheric radius $R_{\rm ph}(t)$ that Zimmerman et al. (2024) calculated in their thorough analysis of SN 2023ixf and give in a table: Supplementary Data (Source Data Table 1). Figure 1 presents the photospheric radius that Zimmerman et al. (2024) calculated at early days (stars and pentagons) as in their Figure 2. We use the same time scale as Zimmerman et al. (2024), who took it from Yaron et al. (2023), but note that the explosion time is uncertain. The time $t=0$ corresponds to the first observed brightening due to the explosion. Yaron et al. (2023) estimated the uncertainty in the time to be about one hour. The explosion at the center of the star has started earlier. In Figure 2, we present the photospheric radii that Zimmerman et al. (2024) calculated at later times.
Figure 1: The time evolution of the photospheric radius from Zimmerman et al. (2024) based on photometry (not on spectroscopy). We fit two straight lines through two groups of points: the first four points, green stars, at $t=1.13-1.42~\rm{day}$ that we mark as Shell 1, and the next 11 points, green pentagons, at $t=1.57-6.13~\rm{day}$ that we mark as Shell 2. The red pentagons are excluded from the fitting; they belong to the phase when the photosphere is moving inward in the mass coordinate of the ejecta. We also indicate the velocity of each shell during the fitting period. The line for Shell 3 fits a later time period, as shown in Figure 2. We identify these shells as photospheric shells appearing one after the other. The straight lines for the three shells are $R_{\rm ph1}=1.050t+0.425$, $R_{\rm ph2}=0.808t+0.477$, and $R_{\rm ph3}=0.445t+0.723$, where the time is in days and the radius in units of $10^{14}~\rm{cm}$. All lines have a coefficient of determination (R-Squared) of $R^{2}=0.98$.
Figure 2: Similar to Figure 1, but including later times, and different scales on the axes. As in 1, red pentagons and pluses are excluded from fittings. We fit a third straight line to the points in the time period of $t=42.19-70.63~\rm{day}$ (green pluses); this is Shell 3.
Zimmerman et al. (2024) argued for a constant photosphere until $t\simeq 2.75~\rm{day}$. We instead find that there are two photospheric shells during that time. We fit two straight lines to the points in Figure 1. Shell 1 that we fit through the first four point (green stars) corresponds to a velocity of $v_{\rm S1}=12,152\pm 4256~\rm{km}~\rm{s}^{-1}$, and Shell 2 that we fit through the green pentagons, to $v_{\rm S2}=9,350\pm 598~\rm{km}~\rm{s}^{-1}$. In Figure 2, we extend the time, and find a time period, represented by the green pluses, where the photosphere increases linearly, this is Shell 3 with a velocity of $v_{\rm S3}=5,152\pm 1214~\rm{km}~\rm{s}^{-1}$. Shell 1 expands, and the photosphere moves outward with it. As it becomes optically thin (transparent), the photosphere expands more slowly as it recedes in the ejecta’s mass coordinate. In the transition from Shell 1 to Shell 2, the photosphere also decreases in radius, suggesting that Shell 1 carries little mass. Then the slower, denser Shell 2 catches up with the photosphere and becomes the photosphere, with linear growth. Later, Shell 2 becomes optically thin, and the photosphere expansion slows down again as it recedes in the ejecta. At about $t\simeq 45~\rm{day}$, Shell 3 catches up and makes the photosphere. The photosphere expands linearly until Shell 3 becomes optically thin at $t\simeq 70~\rm{day}$.
We make the following comments.
(1) From Figure 2, one might propose that there are two shells, where Shell 1 and Shell 2 are part of the same shell. Shell 3 is the same. We find that a straight line fits the first 22 points from $t=1.13~\rm{day}$ to $t=17.81~\rm{day}$. Namely, Shell 1, Shell 2, and the following 7 points form one photospheric shell. The velocity of such a massive early shell is $v_{\rm SE}=7,676\pm 130~\rm{km}~\rm{s}^{-1}$. We consider this less likely (but cannot completely reject this possibility), because (a) the four early points are generally above this line while some of the later points are below it, (b) this line misses some points by more that the error bars, and (c) as we present in Figure 1, Shell 1 and Shell 2 are distinguished.
(2) A shell in this study might be a complete shell covering a solid angle of $\Omega_{\rm sh}=4\pi$ around the explosion site, or a segment of a shell, a cap (or pairs of caps), covering much less, down to $\Omega_{\rm sh}<2\pi$. The cap should be along the line of sight and be large enough to cover a large fraction of the supernova shell towards an observer along its direction. Crudely, a cap should cover an angle of $\gtrsim 0.5\pi$, for which it covers a half of a half sphere to an observer along its axis, to form such a photospheric shell.
(3) Tylenda (2005) found three shells for the luminous red nova V838 Mon. In the case of V838 Mon, the photosphere decreases in radius between the shells. In our analysis, we find a decrease between Shell 1 and Shell 2. This decrease is covered by two observational points (transition from green stars to green pentagons in Figure 1). In the transition from Shell 2 to Shell 3, the photosphere expansion slows down, but its radius does not decrease; the transition occurs over a relatively long time, $t\simeq 6~\rm{day}$ to $t\simeq 45~\rm{day}$.
(4) At late times, a free shell of ejecta expands at a constant velocity (if it does not collide with a massive CSM), as its kinetic energy is much larger than its thermal energy. This is not true very close to the origin, when the ejecta is hot and thermal energy is convected to kinetic energy (adiabatic cooling). Since the stellar radius is $\approx 10^{14}~\rm{cm}$, the velocity of the shells close to the center, $r\lesssim 2\times 10^{14}$, is not constant. A collision with the dense, compact CSM might slow the first shells, while the conversion of thermal to kinetic energy accelerates them near the center. The study of the shell’s behavior close to the star is a subject for a separate study.
3 CCSN remnants with shells
Some CCSNRs present two or three shells, where the shells can be complete (or almost complete), covering a solid angle of $\Omega_{\rm sh}=4\pi$, or be partial, covering a solid angle of $\Omega_{\rm sh}<2\pi$. Such shells, if formed during the explosion, can lead to the formation of two or more photospheric shells during the photospheric phase of the explosion. We present four CCSNRs, all attributed to the JJEM, that show such shell morphologies. They suggest that energetic jet pairs can form such shells, as the JJEM predicts in some cases (but not in all).
In Figure 3 we present two CCSNRs, each with almost full two shells. We mark four points at the boundary of each shell as projected on the plane of the sky with arrows: Shell 1 is the first to form a photosphere, and then Shell 2 catches up with the photosphere of the first shell, and takes over with a slower expansion velocity. Both CCSNRs show clear signatures of at least two pairs of jets. Already Gaensler et al. (1998) argued that SNR G309.2-00.6 was shaped by jets (panel a of Figure 3). Based on its morphology, Soker (2024b) and Shishkin et al. (2024) attributed the explosion of SNR G309.2-00.6 to the JJEM. Soker (2024c) argued that two energetic pairs of jets left their clear marks on CCSNR W44 (panel b of Figure 3). According to the JJEM, more pairs of jets, but with lower energies, participated in the explosion process of W44. The two pairs of jets that Soker (2024c) mark (two double-sided arrows) inflated the two shells that we mark here. This, along with the other images, suggests that two, and possibly three, energetic pairs of jets contributed to the explosion of SN 2023ixf. The JJEM predicts such structures as three-dimensional hydrodynamical simulations show (e.g., Braudo et al. 2025).
Figure 3: Two images of CCSN remnants with two full (or almost full) shells (but not spherical). Studies attributed their morphologies to the JJEM. In each image, we marked the two prominent shells that can form two photospheric shells one after the other during the photospheric phase of such CCSNe; the arrows point to only four points on the projection of each shell boundary onto the plane of the sky. Filled double-lined arrows point at the first shell, and the empty double-lined double-headed arrows point at the second shell. (a) A radio image of SNR G309.2–00.6 adapted from Gaensler et al. (1998), who already argued that jets shaped this CCSN. The rim-nozzle symmetry marks are from Soker (2024b). (b) A composite image of CCSNR W44 from a NASA site. Soker (2024c) added the two double-headed arrows that depict the two axes of the two energetic pairs of jets that participated in the explosion of W44. Magenta: GeV gamma-ray from Fermi’s LAT; yellow: Radio from the Karl G. Jansky Very Large Array; Red: infrared; Blue: X-ray from ROSAT. Credit: NASA/DOE/Fermi LAT Collaboration, NRAO/AUI, JPL-Caltech, ROSAT.
In Figure 4 we present two CCSNRs with partial shells. Namely, they can lead to the observations of two or more photospheric shells only from specific viewing angles. In panel (a), we present SNR G0.9+0.1. Soker (2025a) identified a point symmetric morphology by the structural features marked on the image in pale-blue. We refer to the large ear to the north that has two large front rims. Rims 1 and 2 are the projections of caps, or partial shells, on the plane of the sky. Such rims are formed by jets. If an observer is located along and near the ear direction, the caps of Rim 1 (pointed at by the one-headed yellow arrow) and Rim 2 (double-headed arrow) will form two photospheric shells. The main shell of this CCSNR (pointed at by the triple-headed arrow) will form a third photosphere. In panel (b), we present an image of the CCSNR N63A, with its three pairs of ears that support the JJEM (Soker, 2024d). Ears 2b and 3b are large enough to form large shells at their front (pointed at with the solid yellow arrows). The fronts of these two ears might form a first photospheric shell for an observer facing them. The main shell of N63A (pointed at with the two double-headed arrows) will form the second photospheric shell.
Figure 4: Two images of CCSNRs where at least one shell is partial. (a) A MeerKat radio image at 1.28 GHz of SNR G0.9+0.1 adapted from Heywood et al. (2022). The inset on the upper left is a desaturated image of the pulsar wind nebula. Soker (2025a) added the pale-blue marks of structural features used to identify the point-symmetric morphology. For a line of sight along the large ear in the north, the three shells (their projection on the plane of the sky forms the rims) might form three photospheric shells. The first two (pointed at by the solid one-head yellow arrow and double-lined double-headed yellow arrow) are partial. The third shell (pointed at by the three-headed arrow) is the main shell of the SNR. (b) A Chandra X-ray image of N63A (red, green, blue for different X-ray energy bands); see also Karagöz et al. (2023). Soker (2024d) added the three red lines between the tips of opposite ears to mark the symmetry axis of three pairs of jets that participated in the explosion of this point-symmetric CCSNR. Each ear’s front can form the first photospheric shell, depending on the viewing angle. The main CCSNR forms the next photospheric radius. The light brown region to the upper right of the three red lines is optical light detected by Hubble. (Credit: Enhanced Image by Judy Schmidt based on images provided courtesy of NASA/CXC/SAO & NASA/STScI.)
In addition to these four CCSNe, the Vela CCSNR has two full shells, and SNR G107.7-5.1 has an ear with three rims; the morphologies of both CCSNRs are attributed to the JJEM (Vela: Soker 2023; Soker & Shishkin 2025a; G107.7-5.1: Soker 2024e). About twenty CCSNR morphologies were directly attributed to the JJEM (most others have two dispersed morphologies to be classified). We find six with prominent multiple shell structure. We conclude that jet-shaped morphologies can form two or three, and possibly more, shells during the explosion and lead to the appearance of multiple photospheric shells.
4 Summary
Examining the photospheric time evolution of SN 2023ixf that Zimmerman et al. (2024) calculated, we identify three evolutionary time periods with a constant expansion velocity, as we mark in Figures 1 and 2. We attribute these to three shells of the SN 2023ixf ejecta. Our calculated velocities of these shells during their respective photospheric phases are $v_{\rm S1}\simeq 12,150~\rm{km}~\rm{s}^{-1}$, $v_{\rm S2}\simeq 9,350~\rm{km}~\rm{s}^{-1}$, and $v_{\rm S3}\simeq 5,150~\rm{km}~\rm{s}^{-1}$. The shells can be complete (covering a solid angle of $\Omega_{\rm sh}=4\pi$), or partial if the partial shell, or cap, is along the line of sight.
Several CCSNRs have morphologies with two or more complete or partial shells. We presented four of these in Figures 3 and 4, and mention two more in Section 3. Earlier studies attributed all these CCSNRs to the JJEM, primarily to two or three energetic jet pairs. Additionally, three-dimensional hydrodynamic simulations (Braudo et al., 2025) demonstrate the formation of such structures by consecutive pairs of jets.
Shell 1 of SN 2023ixf becomes optically thin (transparent) very early, at $t=1.4~\rm{day}$, and the photospheric radius decreases (Figure 1). This shell, therefore, contains little mass. This shell may be the front of the ejecta, which has little mass and is not necessarily jet-shaped. However, Shell 2 and Shell 3 are more massive, and, based on CCSNRs that we present in Section 3, we attribute their existence to energetic jets.
We can firmly state that the structure of the three photospheric shells that we identified in SN 2023ixf is fully compatible with jet-shaped CCSNRs and simulations. We take this to support the JJEM. This study adds to the accumulating evidence that the JJEM is the primary explosion mechanism of CCSNE.
Acknowledgements
NS thanks the Charles Wolfson Academic Chair at the Technion for the support.